OpenStudy (anonymous):

How do I simplify (sqrtx+sqrt3)(sqrtx+sqrt27)?

OpenStudy (callisto):

\[(\sqrt x+\sqrt3)(\sqrt x+\sqrt{27})\]\[=\sqrt x(\sqrt x+\sqrt{27})+\sqrt3(\sqrt x+\sqrt{27})\]Can you continue from here?

OpenStudy (anonymous):

Would \[\sqrt{x}\] * \[\sqrt{x}\] cancel out and merely be x?

OpenStudy (callisto):

\[\sqrt{x}\times \sqrt{x} = x\]

OpenStudy (anonymous):

So it would be √x(√x+3√3)+4√3?